Solve the following system of equations. If the system cannot be solved, state that it cannot be solved and explain why.
5x-2y=-24
-4y=-48-10x

Answer by ReadingBoosters(3246)

5x-2y=-24
-4y=-48-10x
...
Put into y=mx+b format
2y =5x+24
y=(5/2)x+12
...
y=(-10/-4)x - (48/-4)
y=(5/2)x + 12
...
They are the same line in a dependent system
where every point on the line is a solution to the system.
-Reading Boosters